Traffic going has a high of 18" and a diameter 7" to the nearest 10th find the amount of space the cone occupies


1
Expert's answer
2021-05-24T18:11:02-0400

"V=\frac1 3Sh=\frac1 3\pi r^2h=\frac1 3 \pi3.5^2*18=230.907056"

Answer: 230.907056